OpenStudy (blackops2_143):

Complete the equation that is related to the given equation. m-18.2=-58 m=-58 + ?

OpenStudy (anonymous):

Idk if it's right but I think it's m=-39.8

OpenStudy (anonymous):

Simplifying M + -18.2 = -58 Reorder the terms: -18.2 + M = -58 Solving -18.2 + M = -58 Solving for variable 'M'. Move all terms containing M to the left, all other terms to the right. Add '18.2' to each side of the equation. -18.2 + 18.2 + M = -58 + 18.2 Combine like terms: -18.2 + 18.2 = 0.0 0.0 + M = -58 + 18.2 M = -58 + 18.2 Combine like terms: -58 + 18.2 = -39.8 M = -39.8 Simplifying M = -39.8 That's how I did it